# µGUIDE: a framework for quantitative imaging via generalized uncertainty-driven inference using deep learning

### v3

@@ 2.1 Framework overview
Figure 2: µGUIDE summarizes information contained in the estimated posterior distributions. A) Examples of degenerate and non-degenerate posterior distributions. B) Presentation of the measures introduced to quantify a posterior distribution on exemplar non-degenerate posterior distributions. Maximum A Posteriori (MAP): is the most likely parameter estimate (dashed vertical lines). Uncertainty: measures the dispersion of the 50% most probable samples using the interquartile range, with respect to the prior range. Ambiguity: measures the Full Width at Half Maximum (FWHM), in percentage with respect to the prior range.
@@ 2.2.1 Comparison with MCMC
Figure 3: Comparison between µGUIDE and MCMC. A) Posterior distributions obtained using either µGUIDE or MCMC on three exemplar simulations with model 2 (SM - $\text{SNR}=50$). B) Bias between the ground truth values used for simulating the diffusion signals, and the maximum-a-posteriori extracted from the posterior distributions using either µGUIDE or MCMC. Sharper and less biased posterior distributions are obtained using µGUIDE.
@@ 3.1 Applicability of µGUIDE to multiple models
Multiple approaches have been used to try and quantify this uncertainty. Gradient descent often provides a measure of confidence for each parameter estimate. Alternative approaches use the shape of the fitted tensor itself as a measure of uncertainty for the fiber direction (Koch et al., 2002; Parker and Alexander, 2003). Other methods also rely on bootstrapping techniques to estimate uncertainty. Repetition bootstrapping for example depends on repeated measurements of signal for each gradient direction, but imply a long acquisition time and cost, and are prone to motion artifacts (Lazar and Alexander, 2005; Jones, 2003). In contrast, residual bootstrapping methods resample the residuals of a regression model. Yet, this approach is heavily dependent on the model and can lead to overfitting (Whitcher et al., 2008; Chung et al., 2006). In general, resampling methods can be problematic for sparse samples, as the bootstrapped samples tend to underestimate the true randomness of the distribution (Kauermann et al., 2009). We propose to quantify the confidence by estimating full posterior distributions, which also has the benefit of highlighting degeneracy.
Variance observed in the posterior distributions can be attributed to several factors. The presence of noise in the signal contributes to irreducible variance, decreasing the confidence in the estimates as the noise level increases. Another source of variance can arise from the choice of acquisition parameters. Different acquisitions may provide varying levels of confidence in the parameter estimates. Under-sampled acquisitions or inadequate b-shells may fail to capture essential information about a tissue microstructure, such as soma or neurite radii, resulting in inaccurate estimates.
@@ 4.1.3 Estimating the likelihood function
$\log\left(p(\boldsymbol{x_{0}}|\boldsymbol{\theta})\right)=-\frac{\sum\left(\boldsymbol{x_{0}}-\sqrt{\mathcal{M}(\boldsymbol{\theta})^{2}+\sigma^{2}}\right)^{2}}{2\sigma^{2}}-m\cdot\log(\sigma\sqrt{2\pi})\enspace,$ (3)
where $\mathcal{M}(\boldsymbol{\theta})$ is the signal obtained using the biophysical model, $\sigma$ is the standard deviation of the Gaussian distributed noise, estimated from the reconstructed magnitude images (Dietrich et al., 2007), and $m$ is the number of observations in the dataset.
@@
The need to compute the signal following the forward model at each iteration makes these sampling methods computationally expensive and time consuming. In addition, they require some adjustments specific to each model, such as the choice of burn-in length, thinning and the number of samples to store. Harms and Roebroeck (2018) recommend to use the Adaptive Metropolis-Within-Gibbs (AMWG) algorithm for sampling dMRI models, initialized with a maximum likelihood estimator obtained from non-linear optimization, with 100 to 200 samples for burn-in and no thinning. The recommended number of samples is model-dependent. Authors recommendations can be found in their paper.
@@ 4.2.1 Neural Posterior Estimator
The conditional probability density function approximators used in this project belong to a class of neural networks called normalizing flows (Papamakarios et al., 2021). These flows are invertible functions capable of transforming vectors generated from a simple base distribution (e.g. the standard multivariate Gaussian distribution) into an approximation of the true posterior distribution. An autoregressive architecture for normalizing flows is employed, implemented via the Masked Autoregressive Flow (MAF) Papamakarios et al. (2017), which is constructed by stacking five Masked Autoencoder for Distribution Estimation models (MADE) Germain et al. (2015). An explanation of how MAF and MADE work is provided in Appendix B.
To test that the predicted posteriors for a given model are not incorrect we use posterior predictive checks, which is described in more details in Appendix C.
@@ 4.3 Application of µGUIDE to biophysical modelling of dMRI data
with $r_{s}$ and $D_{s}$ the soma radius and diffusivity respectively, and $\alpha_{m}$ the $m$th root of $(\alpha r_{s})^{-1}J_{\frac{3}{2}}(\alpha r_{s})=J_{\frac{5}{2}}(\alpha r_{s})$, with $J_{n}(x)$ the Bessel functions of the first kind. We use the six summary statistics defined in (Jallais et al., 2022), which are based on a high and low b-value signal expansion.
@@
To ensure comparability of results, we extract the same number of features $N_{f}$ using the MLP as the number of summary statistics for each model. We therefore use $N_{f}=6$ for the Ball&Stick, the SM and the extended-SANDI models.
@@ 4.3.2 Validation in Numerical Simulations
We start by validating the proposed method using posterior predictive checks (PPC) and simulated signals from model 2 (see more details in Appendix C). Since PPC alone does not guarantee the correctness of the estimated posteriors, we further validated the obtained posterior distributions comparing them with the Adaptive Metropolis-Within-Gibbs MCMC (Roberts and Rosenthal, 2009). We generated simulations following the same acquisition protocol as the real data (see Section 4.3.3), and added Rician distributed noise with a Signal-to-Noise-Ratio (SNR) of 50. We then estimated the posterior distributions using both µGUIDE and the MCMC method implemented in the MDT toolbox (Harms and Roebroeck, 2018). We initialized the sampling using a maximum likelihood estimator. We sampled 15200 samples from the distribution, the first 200 ones being used as burn-in, and no thinning. Similarly, we sampled 15000 samples from the estimated posterior distributions using µGUIDE.
@@ Appendix A Correlation analysis between features extracted by µGUIDE and manually-defined summary statistics
Appendix B Masked Autoregressive Flows
@@
Appendix C Posterior Predictive Checks

### v4

@@ 2.1 Framework overview
Figure 2: µGUIDE summarizes information contained in the estimated posterior distributions. A) Examples of degenerate and non-degenerate posterior distributions. Two Gaussian distributions are fitted to the obtained posterior distribution, where the means and standard deviations are represented by the vertical lines and shaded areas. A voxel is considered as degenerate if the derivative of the fitted Gaussian distributions changes signs more than once (i.e. multiple local maxima), and if the two Gaussian distributions are not overlapping (the distance between the two Gaussian means is inferior to the sum of their standard deviations). B) Presentation of the measures introduced to quantify a posterior distribution on exemplar non-degenerate posterior distributions. Maximum A Posteriori (MAP): is the most likely parameter estimate (dashed vertical lines). Uncertainty: measures the dispersion of the 50% most probable samples using the interquartile range, with respect to the prior range. Ambiguity: measures the Full Width at Half Maximum (FWHM), in percentage with respect to the prior range.
@@ 2.2.1 Comparison with MCMC
Figure 3: Comparison between µGUIDE and MCMC. A) Posterior distributions obtained using either µGUIDE or MCMC on three exemplar simulations with model 2 (SM - $\text{SNR}=50$). Names of the model parameters are indicated in the titles of the panels. B) Bias between the ground truth values used for simulating the diffusion signals, and the maximum-a-posteriori extracted from the posterior distributions using either µGUIDE or MCMC. Sharper and less biased posterior distributions are obtained using µGUIDE.
@@ 3.1 Applicability of µGUIDE to multiple models
Multiple approaches have been used to try and quantify this uncertainty. Gradient descent often provides a measure of confidence for each parameter estimate. Alternative approaches use the shape of the fitted tensor itself as a measure of uncertainty for the fiber direction (Koch et al., 2002; Parker and Alexander, 2003). Other methods also rely on bootstrapping techniques to estimate uncertainty. Repetition bootstrapping for example depends on repeated measurements of signal for each gradient direction, but imply a long acquisition time and cost, and are prone to motion artifacts (Lazar and Alexander, 2005; Jones, 2003). In contrast, residual bootstrapping methods resample the residuals of a regression model. Yet, this approach is heavily dependent on the model and can lead to overfitting (Whitcher et al., 2008; Chung et al., 2006). In general, resampling methods can be problematic for sparse samples, as the bootstrapped samples tend to underestimate the true randomness of the distribution (Kauermann et al., 2009). We propose to quantify the confidence by estimating full posterior distributions, which also has the benefit of highlighting degeneracy. Model-fitting methods with different initializations, as done in e.g. Jelescu et al. (2016), also allow to highlight degeneracies. However, they only provide a partial description of the solution landscape, which can be interpreted as a partial posterior distribution. In contrast, Bayesian methods estimate the full posterior distributions, offering a more accurate and precise characterization of degeneracies and uncertainties. Hence, in this work we decided to use MCMC, a traditional Bayesian method, as benchmark.
Variance observed in the posterior distributions can be attributed to several factors. The presence of noise in the signal contributes to irreducible variance, decreasing the confidence in the estimates as the noise level increases (see Appendix B). Another source of variance can arise from the choice of acquisition parameters. Different acquisitions may provide varying levels of confidence in the parameter estimates. Under-sampled acquisitions or inadequate b-shells may fail to capture essential information about a tissue microstructure, such as soma or neurite radii, resulting in inaccurate estimates.
@@ 4.1.3 Estimating the likelihood function
$\log\left(p(\boldsymbol{x}|\boldsymbol{\theta})\right)=-\frac{\sum\limits_{i=1}^{m}\left(\boldsymbol{x_{i}}-\sqrt{\mathcal{M}(\boldsymbol{\theta})_{i}^{2}+\sigma^{2}}\right)^{2}}{2\sigma^{2}}-m\cdot\log(\sigma\sqrt{2\pi})\enspace,$ (3)
where $\mathcal{M}(\boldsymbol{\theta})$ is the signal obtained using the biophysical model, $\mathcal{M}(\boldsymbol{\theta})_{i}$ is the ith measurement of the signal, $\sigma$ is the standard deviation of the Gaussian distributed noise, estimated from the reconstructed magnitude images (Dietrich et al., 2007), and $m$ is the number of observations in the dataset.
@@
The need to compute the signal following the forward model at each iteration makes these sampling methods computationally expensive and time consuming. In addition, they require some adjustments specific to each model, such as the choice of burn-in length, thinning and the number of samples to store. Harms and Roebroeck (2018) recommend to use the Adaptive Metropolis-Within-Gibbs (AMWG) algorithm for sampling dMRI models, initialized with a maximum likelihood estimator obtained from non-linear optimization, with 100 to 200 samples for burn-in and no thinning. Authors notably investigated the use of starting from the MLE and thinning. They concluded that starting from the MLE allows to start in the stationary distribution of the Markov chain, and has the advantage of removing salt- and pepper-like noise from the resulting mean and standard deviation maps. Their findings also indicate that thinning is unnecessary and inefficient, and they recommend using more samples instead. The recommended number of samples is model-dependent. Authors recommendations can be found in their paper.
@@ 4.2.1 Neural Posterior Estimator
The conditional probability density function approximators used in this project belong to a class of neural networks called normalizing flows (Papamakarios et al., 2021). These flows are invertible functions capable of transforming vectors generated from a simple base distribution (e.g. the standard multivariate Gaussian distribution) into an approximation of the true posterior distribution. An autoregressive architecture for normalizing flows is employed, implemented via the Masked Autoregressive Flow (MAF) Papamakarios et al. (2017), which is constructed by stacking five Masked Autoencoder for Distribution Estimation models (MADE) Germain et al. (2015). An explanation of how MAF and MADE work is provided in Appendix C.
To test that the predicted posteriors for a given model are not incorrect we use posterior predictive checks, which is described in more details in Appendix D.
@@ 4.3 Application of µGUIDE to biophysical modelling of dMRI data
with $r_{s}$ and $D_{s}$ the soma radius and diffusivity respectively, and $\alpha_{m}$ the $m$th root of $(\alpha r_{s})^{-1}J_{\frac{3}{2}}(\alpha r_{s})=J_{\frac{5}{2}}(\alpha r_{s})$, with $J_{n}(x)$ the Bessel functions of the first kind. We use the six summary statistics defined in (Jallais et al., 2022), which are based on a high and low b-value signal expansion. Signal fractions follow the rule $f_{n}+f_{s}+f_{e}=1$, leading to six parameters to estimate for this model.
@@
To ensure comparability of results, we extract the same number of features $N_{f}$ using the MLP as the number of summary statistics for each model. We therefore use $N_{f}=6$ for the Ball&Stick, the SM and the extended-SANDI models. Although the number of features predicted by the MLP is fixed to $N_{f}=6$ for the three models, the characteristics of these six features can be very different, depending on the chosen forward model and the available data (see Appendix A). Training the MLP together with the NPE module allows to maximise inference performance in terms of accuracy and precision.
@@ 4.3.2 Validation in Numerical Simulations
We start by validating the proposed method using posterior predictive checks (PPC) and simulated signals from model 2 (see more details in Appendix D). Since PPC alone does not guarantee the correctness of the estimated posteriors, we further validated the obtained posterior distributions comparing them with the Adaptive Metropolis-Within-Gibbs MCMC (Roberts and Rosenthal, 2009). We generated simulations following the same acquisition protocol as the real data (see Section 4.3.3), and added Gaussian noise to the real and imaginary parts of the simulated signal with a Signal-to-Noise-Ratio (SNR) of 50, and then used the magnitude of this noisy complex signal for our experiments. We then estimated the posterior distributions using both µGUIDE and the MCMC method implemented in the MDT toolbox (Harms and Roebroeck, 2018). We initialized the sampling using a maximum likelihood estimator. We sampled 15200 samples from the distribution, the first 200 ones being used as burn-in, and no thinning. Similarly, we sampled 15000 samples from the estimated posterior distributions using µGUIDE.
@@ Appendix A Correlation analysis between features extracted by µGUIDE and manually-defined summary statistics
Appendix B Impact of noise on the posterior distributions
Noise in the signal impacts the fitting quality of a biophysical model. Fig. 10A shows example posterior distributions for one combination of model 2 parameters, with varying noise levels (no noise, $SNR=50$ and $SNR=25$). Fig. 10B presents uncertainties values obtained on 1000 simulations with varying SNRs. We observe that, as the SNR reduces (i.e. as the noise increases), uncertainty increases. Noise in the signal contributes to irreducible variance. The confidence in the estimates therefore reduces as the noise level increases.
Figure 10: SNR Uncertainty comparison between signals with different noise levels: no noise, $SNR=50$ and $SNR=25$, using model 2. A) Posterior distributions obtained on one example parameter combination (vertical black dashed line) with the three noise levels. B) Histogram of the uncertainty obtained for 1000 signals with different noise levels (in %). Similar ground truths are used for each noise level.
Appendix C Masked Autoregressive Flows
@@
Appendix D Posterior Predictive Checks